Given the following model:

Y= C + I + G + X – Z C = a + bYd
Z = Z0 + zYd
Yd = Y – T

  1. a) Compute the expression for equilibrium income

  2. b) Compute the expression for the tax multiplier

  3. c) Suppose there is an autonomous increase in imports (Z0) of 20 units. To

    counteract this contraction in domestic aggregate demand, assume the government cuts taxes by 20 units. Will equilibrium income rise or fall? By how much? Explain.
Answer #1

(a)

Y = a + b(Y - T) + I + G + X - Z0 - z(Y - T)

Y = a + bY - bT + I + G + X - Z0 - zY + zT

(1 - b + z)Y = a - bT + I + G + X - Z0 + zT

Y = (a - bT + I + G + X - Z0 + zT) / (1 - b + z)

(b)

Tax multiplier = - (b - z) / (1 - b + z) = (z - b) / (1 - b + z)

(c)

Autonomous spending multiplier = 1 / (1 - b + z)

Therefore, Autonomous spending multiplier > Tax multiplier, which means that if autonomous spending rises (falls) and taxes fall (rise) by same amount, there will be a net increase (decrease) in output.

Since imports are leakages from circular flow of income (and they reduce output), an increase in imports is equivalent to a decrease in autonomous spending (like a, I, G or X0).

Net decrease in income = Increase in import (decrease in autonomous spending) = Decrease in tax = 20 units